# Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?

### v3

@@ III-B Criteria for Selection of Codes and Parameters
The eight chosen quantum error correction codes illustrate the historical and conceptual evolution of QECCs. Starting from basic concepts like the quantum repetition code to more complex structures like the heavy-hexagon code, they form a kind of ‘family tree’ that shows the development and diversification of QECCs over time. They encompass a wide array of error correction strategies, from simple error correction (quantum repetition code) to sophisticated techniques like those used in the surface and color codes. This diversity ensures a comprehensive understanding of different error correction methodologies. They vary significantly in terms of complexity and scalability. For example, the Shor code and Steane code represent more traditional approaches, whereas the toric and surface codes are geared toward large-scale quantum computing. This range allows for the assessment of QECCs across different scalability requirements. Different codes are optimized for different error models. Including a variety of codes ensures that the benchmarking study addresses a wide range of error scenarios. Some codes, like the surface code, are designed with practical implementation in mind, considering current quantum computing limitations. This inclusion allows for an assessment of QECCs not just in theory but also in their practical realization. The selection spans codes with varying qubit overheads. For instance, the quantum repetition code is simple but not efficient, whereas codes like the toric code aim to balance error correction efficacy with resource efficiency. These eight codes collectively cover the spectrum of QECCs, from basic to advanced, from theoretical to practical. This range is crucial for a comprehensive benchmarking study. While there are other QECCs, including more might lead to redundancy without significantly adding to the understanding of QECC performance. The chosen codes are sufficiently representative of the broader field. This selection allows for a comparative analysis that can highlight the strengths and weaknesses of different approaches to quantum error correction. It enables a holistic view of the current state of QECC technology.
@@
III-C Approach to Comparative Analysis
The analysis acknowledges that the ‘best’ QECC depends on specific operational contexts, such as the number of qubits available, the type of errors prevalent in the system, and the practical constraints of quantum computing environments. Each selected QECC will be evaluated based on predefined criteria, including qubit overhead, error threshold, error protection, decoding, transversal gates, scalability, realization, and complexity. This structured approach ensures a comprehensive and uniform assessment of each code’s capabilities and limitations. The analysis aims to highlight that while no single QECC is universally superior, each has its strengths and weaknesses. Through this structured and comprehensive approach to comparative analysis, the study will provide valuable insights into the relative merits of different QECCs.
@@ IV-A Comparative Benchmarking of QECCs
Parameters Quantum Repetition Code Shor Code Steane Code Toric Code Surface Code Bacon-Shor Code 3D Color Code Heavy-Hexagon Code Qubit Overhead Error Threshold Error Protection Only half of all bit flip errors. Two qubits, all error type Detects errors on 2 qubits, corrects 1 All Pauli errors All Pauli errors All Pauli errors All Pauli errors All Pauli errors Decoding Automation, ML CNOT/ H Gates CNOT/ H Gates MWPM, Tensor, 10+ Algo MWPM, Tensor, 10+ Algo Statistical Mapping, Check operators Restriction, MaxSAT, 3+ Algo MWPM, ML, Neural Network, 1+ Algo Transversal Gates None None Clifford Gates TPG/Lattice Surgery Lattice Surgery Teleportation Lattice Surgery Lattice Surgery Scalability No No Yes Yes Yes Yes Yes Yes Realization Superconducting, Trapped ion, etc Trapped ion, Optical systems Trapped ion, Rydberg atom Superconducting, Rydberg atom arrays Superconducting, Rydberg atom arrays, Ising anyons Trapped ion None Superconducting Complexity Very Low Low Low Very High High Medium Extremely High Extremely High
Table I shows the parameter analysis for all quantum error correction codes used in this study.
@@
In terms of qubit overhead, it is expressed as a function of the distance variable, . Specifically, the Shor and Steane codes have fixed qubit overheads of and , respectively. The repetition code’s overhead increases linearly with . The overhead for the toric and surface codes grows at a rate proportional to , while for the Bacon-Shor code, it escalates in proportion to . The 3D color code and the heavy-hexagon code exhibit qubit overheads of and , respectively. Figure 9 (left) illustrates the rise in the number of qubits corresponding to increasing values of . For instance, a Bacon-Shor code at a distance demands over qubits, highlighting the significance of qubit overhead as a key benchmarking metric.
@@
The threshold error rate is a critical metric for quantum error correction codes (QECCs), indicating the highest level of physical errors they can effectively correct. The threshold error rates for various codes like the repetition, Shor, Steane, toric, surface, Bacon-Shor, 3D color, and heavy-hexagon codes are [40], [24], [26], [16], [40, 41], [42], [38, 43] and [44] respectively, as depicted in Fig. 9 (right). While the repetition code exhibits the highest threshold, it is important to note its limitation to only detecting bit-flip errors, rendering it impractical for broader applications. Thus, finding a balance in these characteristics is crucial for effective quantum error correction.
@@
Different QECCs have different types and ranges of errors they can correct. The repetition code detects bit flip errors on qubits but does not detect any phase-flip errors [45]. Conversely, the Shor code detects two-qubit errors or corrects an arbitrary single-qubit error [24]. The Steane code is a distance code, detecting errors on two qubits and correcting errors on one qubit [26]. More complex codes like the toric, surface, 3D color, and heavy-hexagon codes can detect and correct all Pauli errors. However, their effectiveness depends on the code’s distance, which is directly related to the qubit overhead and the complexity of implementation. The toric code on a torus is a code, encoding two qubits, while the surface code is a code, encoding only one qubit [46]. The Bacon-Shor code variant, code, has a distance [15]. Unlike the surface code, the color code can suffer from unremovable hook errors due to the specifics of its syndrome extraction circuits, requiring fault-tolerant decoders to use additional flag qubits [47]. In contrast, the heavy-hexagon code is an example of the Bacon-Shor type stabilizer [39].
@@
A decoding algorithm is essential for interpreting stabilizer syndrome values to identify the type and location of errors in a quantum error correction code (QECC). These algorithms introduce additional complexity to QECCs, and the performance of these codes is highly dependent on such algorithms. The availability of multiple decoding algorithms enhances the versatility of a code. The repetition code can use classical 2D Ising model decoders [48] or machine learning-based approaches [49]. The Shor and Steane codes employ CNOT and Hadamard gates for decoding [50]. The toric and surface codes are supported by a variety of decoding algorithms, including minimum weight perfect-matching (MWPM) [51], tensor decoders [52], union-find decoders [53], brief perfect matching [54], renormalization group methods [55], Markov-chain Monte Carlo techniques [56], cellular automata [57], neural networks [58], sliding and parallel window decoders [59], generalized belief propagation [60], among others. The Bacon-Shor code uses the mapping effect of noise to statistical mechanical models for decoding [61]. Although its check operators are few-body, stabilizer weights scale with the number of qubits, and stabilizer expectation values are obtained by taking products of gauge-operator expectation values [62]. The 3D color code may utilize various decoders such as the projection decoder [63], integer-program-based decoder [64], restriction decoder [38], cellular-automation decoder [65], or the MaxSAT-based decoder [66]. The heavy-hexagon code has two decoders available primarily for itself, one using machine learning [44] and the other using a neural network [67], along with MWPM.
@@
Transversal gates facilitate operations on logically encoded qubits without the need for decoding. The choice of transversal gate contributes to the complexity of implementing a QECC. The repetition and Shor codes do not rely on transversal gates and are infrequently scalable across multiple qubits. The Steane code employs single-qubit Clifford gates, realizing the binary octahedral subgroup [68]. The toric, surface, 3D color, and heavy-hexagon codes utilize lattice surgery [69]. In the toric code, transversal Pauli gates can be applied to non-trivial loops [70]. In symmetric Bacon-Shor codes, the Logical Hadamard is transversal up to a qubit permutation and can be implemented via teleportation [71].
@@
The practical implementations of QECCs reflect their applicability in real-world scenarios, and the range of qubit technologies they can be implemented on indicates their versatility. The repetition code has been realized on superconducting qubits [72, 73], trapped-ion qubits [74], liquid-state NMR [75], and NV diamond qubits [76]. The Shor code has seen implementation on trapped-ion qubits [77] and in optical systems [78]. The Steane code has been realized with trapped-ion qubits [79] and Rydberg atom arrays [80]. Both the toric and surface codes have been implemented using Ising anyons [81], while the surface code has also been realized with superconducting qubits [82] and Rydberg atomic arrays [80]. The Bacon-Shor code’s realization has been limited to trapped-ion qubits [83]. As of now, there has been no practical realization of the 3D color code. The heavy-hexagon code has been implemented in superconducting qubits [84].
@@
The repetition and Shor codes are not scalable across multiple qubits, in contrast to the Steane, toric, surface, Bacon-Shor, 3D color, and heavy-hexagon codes, which are scalable. The implementation complexity of these codes can be gauged from their circuit or lattice structures. The repetition code has a very low complexity. The Shor and Steane codes can be categorized as having low complexity. The surface code, with its square lattice structure, is considered highly complex. Similarly, the toric code, despite its resemblance in structure, is deemed highly complex due to its encoding of two qubits instead of one. The Bacon-Shor code falls into the medium complexity category. Finally, the lattice structures of the 3D color and heavy-hexagon codes indicate their extremely high complexity in terms of implementation.
@@
Now that we have analyzed QECCs in terms of their parameters, we can collectively benchmark them using a radar chart as shown in Fig. 10. This chart consists of eight axes, each representing a different parameter. Colored lines connecting the axes represent the QECCs. The ‘qubit overhead’ parameter in the chart varies from constants to the highest order . The ‘error threshold’ spans from the lowest to the highest order value. ‘Error protection’ encompasses types ranging from basic bit-flip to all Pauli errors. ‘Decoding’ varies from a single classical decoder to the maximum number available for any QECC. ‘Transversal gates’ extend from none to the most complex, lattice surgery. ‘Scalability’ is binary, indicated by yes or no. ‘Realization’ varies from none to the maximum number of qubit technologies on which a QECC has been implemented. The ‘complexity’ of a QECC is rated on a scale of six levels, from very low to extremely high.
@@
Considering the surface codes and comparing them with the radar chart in Fig. 10, we observe that they offer the highest form of error protection with scalability and have been realized on a considerable number of qubit technologies, with multiple decoders available. Despite having a relatively good error threshold, making them a viable option, they come at the cost of high qubit overhead, complex implementation, and the need to utilize the most complex transversal gate. Conversely, the repetition code features the least complex transversal gate and is straightforward to implement, but offers limited error protection. The choice of QECC thus becomes an art of balancing parameters and requirements.
@@ IV-B Benchmarking QECC Variations
V How to Strategically Choose a QECC?
@@ V-A An Universal Streamlined Approach
The process of selecting a code is dynamic and situational, with no fixed rules. Hence, paying attention to minor differences in codes is as important as considering major codes. In terms of code selection, although a universal solution doesn’t exist, we can establish a general systematic strategy based on typical priorities. The selection strategy for QECCs, outlined in Fig. 13, involves a multi-stage filtration process, starting with an initial evaluation based on error types, qubit type, and feasibility. This is followed by a detailed examination of error thresholds and qubit overhead. The strategy emphasizes optimizing qubit overhead and scalability, revisiting previous stages if overhead is excessive or scalability is lacking. Complexity and transversal gates are then assessed to minimize both, while maximizing decoding algorithm options. The final QECC choice balances complexity against decoding efficiency, tailored to the scenario’s demands. A last review ensures the chosen QECC meets all criteria, with adjustments made to internal parameters as needed. If inadequacies are found, the process recommences.
V-B Introducing Our Tool for Strategic Recommendation

### v4

@@ III-B Criteria for Selection of Codes and Parameters
The nine chosen quantum error correction codes illustrate the historical and conceptual evolution of QECCs. Starting from basic concepts like the quantum repetition code to more complex structures like the heavy-hexagon code, they form a kind of ‘family tree’ that shows the development and diversification of QECCs over time. They encompass a wide array of error correction strategies, from simple error correction (quantum repetition code) to sophisticated techniques like those used in the surface and color codes. This diversity ensures a comprehensive understanding of different error correction methodologies. They vary significantly in terms of complexity and scalability. For example, the Shor code and Steane code represent more traditional approaches, whereas the toric and surface codes are geared toward large-scale quantum computing. This range allows for the assessment of QECCs across different scalability requirements. Different codes are optimized for different error models. Including a variety of codes ensures that the benchmarking study addresses a wide range of error scenarios. Some codes, like the surface code, are designed with practical implementation in mind, considering current quantum computing limitations. This inclusion allows for an assessment of QECCs not just in theory but also in their practical realization. The selection spans codes with varying qubit overheads. For instance, the quantum repetition code is simple but not efficient, whereas codes like the toric code aim to balance error correction efficacy with resource efficiency. These nine codes collectively cover the spectrum of QECCs, from basic to advanced, from theoretical to practical. This range is crucial for a comprehensive benchmarking study. While there are other QECCs, including more might lead to redundancy without significantly adding to the understanding of QECC performance. The chosen codes are sufficiently representative of the broader field. This selection allows for a comparative analysis that can highlight the strengths and weaknesses of different approaches to quantum error correction. It enables a holistic view of the current state of QECC technology.
@@
III-C Approach for Comparative Analysis
The analysis acknowledges that the ‘best’ QECC depends on specific operational contexts, such as the number of qubits available, the type of errors prevalent in the system, and the practical constraints of quantum computing environments. Each selected QECC will be evaluated based on predefined criteria, including qubit overhead, error threshold, error protection, decoding, transversal gates, scalability, realization, and complexity. This structured approach ensures a comprehensive and uniform assessment of each code’s capabilities and limitations. The analysis aims to highlight that while no single QECC is universally superior, each has its strengths and weaknesses. Through this structured and comprehensive approach to comparative analysis, the study will provide valuable insights into the relative merits of different QECCs. Table I shows the parameter analysis for all quantum error correction codes used in this study.
@@ IV-A Comparative Benchmarking of QECCs
Parameters Quantum Repetition Code Shor Code Steane Code Toric Code Surface Code Bacon-Shor Code 3D Color Code Heavy-Hexagon Code BB (or Gross) Code Qubit Overhead Error Threshold Error Protection Only half of all bit flip errors. Two qubits, all error type Detects errors on 2 qubits, corrects 1 Pauli errors Pauli errors Pauli errors Pauli errors Pauli errors Pauli errors Decoding Automation, ML CNOT/ H Gates CNOT/ H Gates MWPM, Tensor, 10+ Algo MWPM, Tensor, 10+ Algo Statistical Mapping, Check operators Restriction, MaxSAT, 3+ Algo MWPM, ML, Neural Network, 1+ Algo BP-OSD Transversal Gates None None Clifford Gates TPG/Lattice Surgery Lattice Surgery Teleportation Lattice Surgery Lattice Surgery Lattice Surgery Scalability No No Yes Yes Yes Yes Yes Yes Yes Realization Superconducting, Trapped ion, etc Trapped ion, Optical systems Trapped ion, Rydberg atom Superconducting, Rydberg atom arrays Superconducting, Rydberg atom, Ising anyons Trapped ion None Superconducting None Complexity Very Low Low Low Very High High Medium Extremely High Extremely High Very High
@@
Qubit overhead is expressed as a function of the distance variable, . Specifically, the Shor and Steane codes have fixed qubit overheads of and , respectively. The repetition code’s overhead increases linearly with . The overhead for the toric and surface codes grows at a rate proportional to , while for the Bacon-Shor code, it escalates in proportion to . The 3D color code and the heavy-hexagon code exhibit qubit overheads of and , respectively. The BB codes belong to a wider category of codes characterized by their low overhead. To quantify this overhead, it can be expressed as . where , indicating a subquadratic relationship. The specific values of and vary based on the particular BB code constructions and might not be consistent across different versions. Thus, for our analysis, we will concentrate on a specific instance known as the Gross Code. The overhead for Gross Codes increases nearly linearly, at a factor of times the distance . Figure 10 (left) illustrates the rise in the number of qubits corresponding to increasing values of . For instance, a Bacon-Shor code at a distance demands over qubits, highlighting the significance of qubit overhead as a key benchmarking metric.
@@
The threshold error rates for various codes like the repetition, Shor, Steane, toric, surface, Bacon-Shor, 3D color, heavy-hexagon and BB codes are [43], [26], [28], [18], [43, 44], [45], [40, 46], [47] and [42] respectively, as depicted in Fig. 10 (right). While the repetition code exhibits the highest threshold, it is important to note its limitation to only detecting bit-flip errors, rendering it impractical for broader applications. Thus, finding a balance in these characteristics is crucial for effective quantum error correction.
@@
Different QECCs have different types and ranges of errors they can correct. The repetition code detects bit flip errors on qubits but does not detect any phase-flip errors [48]. Conversely, the Shor code detects two-qubit errors or corrects an arbitrary single-qubit error [26]. The Steane code is a distance code, detecting errors on two qubits and correcting errors on one qubit [28]. More complex codes like the toric, surface, 3D color, heavy-hexagon and BB codes can detect and correct all Pauli errors. However, their effectiveness depends on the code’s distance, which is directly related to the qubit overhead and the complexity of implementation. The toric code on a torus is a code, encoding two qubits, while the surface code is a code, encoding only one qubit [49]. The Bacon-Shor code variant, code, has a distance [17]. Unlike the surface code, the color code can suffer from unremovable hook errors due to the specifics of its syndrome extraction circuits, requiring fault-tolerant decoders to use additional flag qubits [50]. In contrast, the HH code is an example of the Bacon-Shor type stabilizer [41].
@@
A decoding algorithm is essential for interpreting stabilizer syndrome values to identify the type and location of errors in a quantum error correction code (QECC). These algorithms introduce additional complexity to QECCs, and the performance of these codes is highly dependent on such algorithms. The availability of multiple decoding algorithms enhances the versatility of a code. The repetition code can use classical 2D Ising model decoders [51] or machine learning-based approaches [52]. The Shor and Steane codes employ CNOT and Hadamard gates for decoding [53]. The toric and surface codes are supported by a variety of decoding algorithms, including minimum weight perfect-matching (MWPM) [54], tensor decoders [55], union-find decoders [56], brief perfect matching [57], renormalization group methods [58], Markov-chain Monte Carlo techniques [59], cellular automata [60], neural networks [61], sliding and parallel window decoders [62], generalized belief propagation [63], among others. The Bacon-Shor code uses the mapping effect of noise to statistical mechanical models for decoding [64]. Although its check operators are few-body, stabilizer weights scale with the number of qubits, and stabilizer expectation values are obtained by taking products of gauge-operator expectation values [65]. The 3D color code may utilize various decoders such as the projection decoder [66], integer-program-based decoder [67], restriction decoder [40], cellular-automation decoder [68], or the MaxSAT-based decoder [69]. The heavy-hexagon code has two decoders available primarily for itself, one using machine learning [47] and the other using a neural network [70], along with MWPM. The BB code can be decoded using an extended version of the BP-OSD decoder [71] to account for measurement errors.
@@
Transversal gates facilitate operations on logically encoded qubits without the need for decoding. The choice of transversal gate contributes to the complexity of implementing a QECC. The repetition and Shor codes do not rely on transversal gates and are infrequently scalable across multiple qubits. The Steane code employs single-qubit Clifford gates, realizing the binary octahedral subgroup [72]. The toric, surface, 3D color, heavy-hexagon and BB codes utilize lattice surgery [73]. In the toric code, transversal Pauli gates can be applied to non-trivial loops [74]. In symmetric Bacon-Shor codes, the Logical Hadamard is transversal up to a qubit permutation and can be implemented via teleportation [75].
@@
The practical implementations of QECCs reflect their applicability in real-world scenarios, and the range of qubit technologies they can be implemented on indicates their versatility. The repetition code has been realized on superconducting qubits [76, 77], trapped-ion qubits [78], liquid-state NMR [79], and NV diamond qubits [80]. The Shor code has seen implementation on trapped-ion qubits [81] and in optical systems [82]. The Steane code has been realized with trapped-ion qubits [83] and Rydberg atom arrays [84]. Both the toric and surface codes have been implemented using Ising anyons [85], while the surface code has also been realized with superconducting qubits [86] and Rydberg atomic arrays [84]. The Bacon-Shor code’s realization has been limited to trapped-ion qubits [87]. The heavy-hexagon code has been implemented in superconducting qubits [88]. As of now, there has been no practical realization of the 3D color and the BB codes.
@@
The repetition and Shor codes are not scalable across multiple qubits, in contrast to the Steane, toric, surface, Bacon-Shor, 3D color, heavy-hexagon, and BB codes, which are scalable. The implementation complexity of these codes can be gauged from their circuit or lattice structures. The repetition code has a very low complexity. The Shor and Steane codes can be categorized as having low complexity. The surface code, with its square lattice structure, is considered highly complex. Similarly, the toric code, despite its resemblance in structure, is deemed highly complex due to its encoding of two qubits instead of one. Despite its low qubit overhead, the Gross Code is also considered highly complex due to its capability to encode multiple qubits simultaneously and its requirement for a two-layered setup. The Bacon-Shor code falls into the medium complexity category. Finally, the lattice structures of the 3D color and heavy-hexagon codes indicate their extremely high complexity in terms of implementation.
@@
Fig. 11 benchmarks the QECCs using a radar chart which consists of eight axes, each representing a different parameter. Colored lines connecting the axes represent the QECCs. The ‘qubit overhead’ parameter in the chart varies from constants to the highest order . The ‘error threshold’ spans from the lowest to the highest order value. ‘Error protection’ encompasses types ranging from basic bit-flip to all Pauli errors. ‘Decoding’ varies from a single classical decoder to the maximum number available for any QECC. ‘Transversal gates’ extend from none to the most complex, lattice surgery. ‘Scalability’ is binary, indicated by yes or no. ‘Realization’ varies from none to the maximum number of qubit technologies on which a QECC has been implemented. The ‘complexity’ of a QECC is on a scale of six levels, from very low to extremely high.
@@
We observe that surface codes offer the highest form of error protection with scalability and have been realized on a considerable number of qubit technologies, with multiple decoders available. Despite having a relatively good error threshold, making them a viable option, they come at the cost of high qubit overhead, complex implementation, and the need to utilize the most complex transversal gate. Conversely, the repetition code features the least complex transversal gate and is straightforward to implement, but offers limited error protection. The choice of QECC thus becomes an art of balancing parameters and requirements.
@@ IV-B Benchmarking QECC Variations
V Strategic Choice of a QECC
@@ V-A An Universal Streamlined Approach
The process of selecting a code is dynamic and situational, with no fixed rules. Hence, paying attention to subtle differences in codes is as important as considering major codes. In terms of code selection, although a universal solution doesn’t exist, we can establish a general systematic strategy based on typical priorities. The selection strategy for QECCs, outlined in Fig. 14, involves a multi-stage filtration process. The strategic ordering of the multi-stage filtration process in selecting a QECC is designed to ensure a systematic and efficient narrowing down of potential codes to the one most suited for a specific scenario. This process begins with the initial evaluation of error types, qubit type, and feasibility, prioritizing the fundamental compatibility with the physical qubit system and the types of errors anticipated. This foundational step ensures that any QECC considered has the basic capability to address the primary challenges of the quantum computing system in question. Following this, the process assesses practicality through an examination of error thresholds and qubit overhead, which are critical for determining the feasibility of implementing a QECC in real-world applications. Addressing these aspects early in the selection process ensures that codes requiring unrealistic additional resources or that cannot manage expected error rates are filtered out before considering more nuanced factors. Scalability is addressed subsequently to ensure that the chosen code not only meets the immediate functional and practical requirements but also holds potential for future expansion and adaptation. This prioritization ensures that scalability does not compromise essential efficacy and practicality. Finally, the selection process considers the complexity of implementing the QECC, the ability to perform transversal gates, and the balance between complexity and decoding efficiency. These later stages refine the selection from among codes that have already passed the more fundamental tests, focusing on optimizing the balance between practical implementation and efficient operation. By following this logical sequence, the selection process ensures that the chosen QECC is not only theoretically capable of correcting anticipated errors but is also practically implementable, scalable, and efficient, balancing immediate needs with long-term potential. A last review ensures the chosen QECC meets all criteria, with adjustments made to internal parameters as needed. Identifying inadequacies in a QECC upon final review—such as poor error correction, high qubit overhead, scalability issues, or implementation complexities—necessitates restarting the selection process. This approach enables iterative refinement, using initial insights to more accurately meet scenario needs, accommodates evolving considerations, and adjusts to changes in resources, technology, or objectives. Thus, reassessment ensures the final QECC choice remains optimal.
V-B Strategic Recommendation Tool